Use either elimination or substitution to find the solutions (if any)
Graph the two lines, labeling the x-intercepts, y-intercepts, and points of intersection


y = 2x + 3 and y = -x – 4

Answer by mananth(16949)

y=2x+3
y=-x-4
equation1
when x= 0 y= 3 ( 0 , 3 )
when y = 0 x= -1.5 ( -1.5 , 0 )
Equation 2
when x= 0 y= -4 ( 0 , -4 )
when y = 0 x= -4 ( -4 , 0 )